OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (anonymous):

I know the other zero is 1 - 3i but I don't remember how to do this...

OpenStudy (anonymous):

... what are the zeros of (x-a)(x-b)(x-c) ta da ;D you got this

OpenStudy (anonymous):

\[(x-2)(x+4)(x-(1+3i))(x-(1-3i))\]

OpenStudy (anonymous):

the last pair is not hard, you can work backwards if the zero is \(1+3i\) then you can write \[x=1+3i\] \[x-1=3i\] \[(x-1)^2=-9\] \[x^2+2x+1=-9\] \[x^2+2x+10=0\]

OpenStudy (anonymous):

or you can remember that if the zero is \(a+bi\) the quadratic is \[x^2-2ax+a^2+b^2\] either way

OpenStudy (anonymous):

i made a typo above, it is \[x^2-2x+10\]

OpenStudy (anonymous):

x^4 - 2x^2 + 36x - 80 ?

OpenStudy (anonymous):

yesh @allison_ good job